HUBBLE SPACE TELESCOPE DAILY REPORT #1381 PERIOD COVERED: 0000Z (UTC) 05/01/95 - 0000Z (UTC) 05/02/95 Daily Status Report as of 122/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30-minute long, dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxies 3C225.0 and 3C61.1 as part of a continuum survey and to study the galaxies' morphology. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 5772 (PC Imaging of Gravitational Lenes, WF/PC-2 Augmentation, Carryover) The WF/PC-2 was used to make deep exposure observations of the gravitational lens system 1120+019. The images were obtained in two band passes, with the longer wavelength filter intended for the lensing galaxy and the shorter for the quasar. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 6142 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Smaller Set Part II) The WF/PC-2 was used to make Earth Flat calibration images for the ultraviolet and narrow-band filters to correct for the telescope's illumination pattern. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FOS/RD 5441 (The Inner Regions of Quasars: Cycle 4) The Faint Object Spectrograph (Red Side) was used to make spectrographic observations of the variable quasar TEX1156+21. The observations will be correlated with radio and X-Ray data for this object. There was an FGS #2 guide star acquisition failure to coarse track at 121/0752Z (HSTAR 5195, Reported yesterday). The scan step limit was exceeded, and FGS #2 was dropped from the acquisition logic and FGS #1 failed to coarse track backup. Total coarse track time was less than 50 minutes. No other anomalies were noted. 1.6 Completed WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. The observations were taken in parallel with other prime science observations. The observations were completed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 4819 (Point Spread Function Characterization) The WF/PC-2 was used to make multiple exposures of the open cluster M-67 in Cancer (aka NGC-2682) during coarse and fine dwell scans to study the point spread function. A Dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed GHRS/1 5712 (Local Interstellar Medium and D/H Ratio) The High Resolution Spectrograph (Side One) was used to make observations of the standard star HD128620 (aka, Alpha Cent) to study the local interstellar medium which lies along the line of sight to the target star. The goal is to observe interstellar H and D in absorption against the stellar emission lines, and thereby determine the H to D ratio. The observations were completed as planned, and no anomalies were noted. 1.9 Completed Two Sets of WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make both dark and visual flats. The observations were completed as planned, and no problems were reported. 1.10 Completed WF/PC-2 5663 (WF/PC-2 Cycle 4: Photometric Calibration Monitor: Fields Part 2) The WF/PC-2 was used to make observations of globular cluster NGC-5139 with various filters in order to check the photometric calibration of the WF/PC-2. The observations were completed as planned, and no anomalies were reported. 1.11 Completed GHRS/2 5671 (GHRS Dark Noise Monitor: Cycle 4) The High Resolution Spectrograph (Side Two) was used to make dark calibration images. These were strictly internal tests for routine monitoring of detector dark noise. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 5454 (Ellipticals with Kinematically Distinct Cores: Cycle 4 Medium) The WF/PC-2 was used to make observations of the small elliptical galaxy NGC-4552 which contains a massive core which is kinematically distinct from the rest of the galaxy. FGS #3 had a guide star acquisition failure at 121/2030Z (HSTAR 5197). A search radius limit was exceeded. Preceding the acquisition there were two roll maneuver updates and a T2 slew. Following this was a roll delay, a slew, and two full maneuver updates, all of which were successful. 1.13 Completed FOS/RD/S/C 5528 (FOS Cycle 4: Focus, X-Patch, Y-Patch) The Faint Object Spectrograph (Both the Red and Blue Sides) was used to make both Dark and WAVE calibration observations. No external sources were used. The observations were completed as planned, and no anomalies were reported. 1.14 Completed GHRS/1 5594 (GHRS Side 1 Sensitivity Monitor: Cycle 4) The High Resolution Spectrograph (Side One) was used to make calibration observations of star Mu-Col to monitor the long-term sensitivity of the system. The GHRS/1 Special Lamp #2 was reported to show high current for one observation at 122/0106Z (HSTAR 5196). No other anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 6 At 121/2030Z (See Note 1.12) Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 17 Successful : 16 A full maneuver update on trackers #1 and 2 failed at 121/1257Z. The next maneuvers were okay. The problem is under investigation. 2.3 Significant HSTARs: None 2.4 Operations Notes: No operational problems were reported. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-